(D2 - 4 ) y = xex

1
Expert's answer
2021-05-26T22:56:32-0400

Ans:-

"(D^2 - 4 ) y = xe^x"

The Auxiliary equation will be "\ m^2-4=0, \ m=2,2\\"

"C.F.=(Ax+B)e^{2x}"

"P.I.=\dfrac{1}{D^2-4}xe^{x}\\"


"=4e^x (-1+\dfrac{D^2}{4})^{-1}\times x"


"=""\ \ \ 4e^x"

Hence the complete solution is

"Y=(Ax+B)e^{2x}+4e^x"